# Diffusion Models for Intelligent Transportation Systems: A Survey

### v1

@@ III-B Advantages of Diffusion Models
TABLE I: Applications of diffusion models in intelligent transportation systems. To classify existing models, three key criteria are considered: the task, the denoising condition, and the architecture. Additionally, the datasets and open source are provided. The following abbreviations are used to denote the architectures: DDPM (Denoising Diffusion Probabilistic Model) [16], DDIM (Denoising Diffusion Implicit Model) [73], ADM (Ablated Diffusion Model) [45], LDM (Latent Diffusion Model) [38], LED (LEapfrog Diffusion Model) [74], VLDM (Video Latent Diffusion Model) [57], EDM (Elucidating Diffusion Model) [75], FDM (Flexible Diffusion Model) [76], D3PM (Discrete Denoising Diffusion Probabilistic Model) [77], CARD (Classification and Regression Diffusion Model) [78].
@@
DiffusionDet [79] 2D object detection conditioned on image feature DDIM CrowdHuman [80]
COCO [81]
@@
DetDiffusion [82] 2D object detection conditioned on perception-aware attributes LDM COCO [81] 2024 CVPR ——
@@
conditioned on BEV feature DDPM nuScenes [84] 2024 AAAI DiffBEV
@@
conditioned on image feature DDIM ADE20K [85]
NYU-DepthV2 [86]
KITTI [87] et al.
@@
conditioned on text LDM ADE20K [85]
RefCOCO [89]
NYU-DepthV2 [86]
@@
Chen et al. [90] multi-object tracking conditioned on text LDM MOT20 [91] et al. 2024 CVPR LtD-MOT
Luo et al. [92] multi-object tracking conditioned on two adjacent raw images DDPM MOT20 [91] et al. 2024 AAAI DiffusionTrack
Xie et al. [93] object tracking unconditional DDIM GOT-10k [94]
LaSOT [95]
@@
Luo et al. [96] 3D point cloud generation conditioned on shape latent [97] DDPM ShapeNet 2021 CVPR DPC
@@
conditioned on text LDM VOC [98]
ADE20K [85]
Cityscapes [99]
@@
DatasetDM [40] perception data augmentation conditioned on text LDM COCO [81] et al. 2023 NIPS DatasetDM
MID [27] human trajectory prediction conditioned on observed trajectories DDPM SDD [100]
ETH [101]
UCY [102]
@@
conditioned on observed trajectories LED SDD [100] et al. 2023 CVPR LED
@@
conditioned on observed scene DDIM ETH [101] et al. 2024 CVPR SingularTrajectory
@@
conditioned on observed trajectories
conditioned on endpoint
DDPM SDD [100] et al. 2024 arxiv ——
@@
conditioned on gate DDPM SDD [100] et al. 2024 NIPS ——
MotionDiffuser [28] multi-agent prediction conditioned on observed scene , constraints
@@
LDM WOMD [107] 2023 CVPR ——
SceneDiffusion [58] multi-agent prediction conditioned on observed scene, interval time
@@
LDM Argoverse [108] 2023 ITSC ——
Equidiff [29] vehicle trajectory prediction conditioned on observed trajectories, interactions DDPM NGSIM [109] 2023 ITSC ——
Yao et al. [110] vehicle trajectory prediction conditioned on observed trajectories, map DDPM Argoverse2 [111] 2023 CSIS-IAC ——
Diffuser [46] behavior planning unconditional
@@
conditioned on rewards, constraints, skills
@@
ADM D4RL [112] 2023 ICLR ——
MPD [47] motion planning unconditional
@@
Diffusion-ES [113] motion planning unconditional truncated DDPM nuPlan [114] 2024 CVPR diffusion-es
@@
conditioned on adjacent views VLDM nuScenes [84] 2024 CVPR Drive-WM
@@
conditioned on past frame, text VLDM WOMD [107] et al. 2024 CVPR DriveAGI
CTG [48] vehicle trajectory generation conditioned on observed scene
@@
ADM nuScenes [84] 2023 ICRA CTG
CTG++ [49] multi-agent trajectory generation conditioned on observed scene
@@
ADM nuScenes [84] 2023 CoRL CTG++
Dragtraffic [59] multi-agent trajectory generation conditioned on initial scene, text LED WOMD [107] 2024 IROS Dragtraffic
DJINN [50] multi-agent trajectory generation conditioned on arbitrary state
classifier-free guidance
@@
Argoverse2 [111] 2023 NIPS ——
Rempe et al. [53] human trajectory generation conditioned on observed scene
@@
FDM [76] image-based driving scenario generation conditioned on previously sampled frames FDM Carla [116] 2022 NIPS ——
GAIA-1 [54] image-based driving scenario generation conditioned on past image, text, action tokens
@@
DriveDreamer-2 [117] image-based driving scenario generation conditioned on structured info by LLMs, text EDM nuScenes [84] 2024 arxiv DriveDreamer2
@@
nuScenes [84] 2024 CVPR panacea
@@
nuScenes [84] 2023 arxiv DrivingDiffusion
WoVoGen [63] image-based driving scenario generation conditioned on past world volumes, actions, text, 2D image feature LDM nuScenes [84] 2023 arxiv WoVoGen
LiDMs [64] point cloud-based driving scenario generation unconditional
@@
Copilot4D [55] point cloud-based driving scenario generation conditioned on past observations, actions
@@
DiffTraj [34] GPS trajectory generation conditioned on trip region, departure time et al.
@@
DiffSTG [36] traffic flow forecasting Conditioned on past graph signals, graph structure DDIM PEMS [121] et al. 2023 GIS DiffSTG
@@
conditioned on past graph signals feature, adjacency matrix DDPM PEMS [121] et al. 2024 arxiv SpecSTG
@@
ST-SSPD [124] traffic flow forecasting conditioned on past data points, temporal encoding, node identifier DDPM METR-LA [125] et al. 2023 MobiArch ——
@@
DiffTAD [127] trajectory anomaly detection unconditional DDIM NGSIM [109] 2024 KBS ——
VAD [128] video anomaly detection unconditional
@@
LDM, DDIM CUHK Avenue [129] et al. 2023 ICCV ——

### v3

@@ Appendix A Supplementary Materials
TABLE I: Applications of diffusion models in intelligent transportation systems. Three key criteria are considered to classify existing models: the task, the denoising condition, and the architecture. Additionally, the datasets and open source are provided. The following abbreviations are used to denote the architectures: DDPM (Denoising Diffusion Probabilistic Model) [16], DDIM (Denoising Diffusion Implicit Model) [158], ADM (Ablated Diffusion Model) [34], LDM (Latent Diffusion Model) [32], LED (LEapfrog Diffusion Model) [87], VLDM (Video Latent Diffusion Model) [49], EDM (Elucidating Diffusion Model) [162], FDM (Flexible Diffusion Model) [121], D3PM (Discrete Denoising Diffusion Probabilistic Model) [130], CARD (Classification and Regression Diffusion Model) [163].
@@
DiffusionDet [75] 2D object detection conditioned on image feature DDIM CrowdHuman
COCO
@@
DetDiffusion [76] 2D object detection conditioned on perception-aware attributes LDM COCO 2024 CVPR ——
@@
conditioned on BEV feature DDPM nuScenes 2024 AAAI DiffBEV
@@
conditioned on image feature DDIM ADE20K
NYU-DepthV2
KITTI et al.
@@
conditioned on text LDM ADE20K
RefCOCO
NYU-DepthV2
@@
Chen et al. [79] multi-object tracking conditioned on text LDM MOT20 et al. 2024 CVPR LtD-MOT
Luo et al. [80] multi-object tracking conditioned on two adjacent raw images DDPM MOT20 et al. 2024 AAAI DiffusionTrack
Xie et al. [81] object tracking unconditional DDIM GOT-10k
LaSOT
@@
Luo et al. [83] 3D point cloud generation conditioned on shape latent DDPM ShapeNet 2021 CVPR DPC
@@
conditioned on text LDM VOC
ADE20K
Cityscapes
@@
DatasetDM [82] perception data augmentation conditioned on text LDM COCO et al. 2023 NIPS DatasetDM
MID [28] human trajectory prediction conditioned on observed trajectories DDPM SDD
ETH
UCY
@@
conditioned on observed trajectories LED SDD et al. 2023 CVPR LED
@@
conditioned on observed scene DDIM ETH et al. 2024 CVPR SingularTrajectory
@@
conditioned on observed trajectories, endpoint DDPM SDD et al. 2024 arxiv ——
@@
conditioned on gate DDPM SDD et al. 2024 NIPS ——
MotionDiffuser [29] multi-agent prediction conditioned on observed scene, constraints;
@@
LDM WOMD 2023 CVPR ——
SceneDiffusion [50] multi-agent prediction conditioned on observed scene, interval time;
@@
LDM Argoverse 2023 ITSC ——
Equidiff [30] vehicle trajectory prediction conditioned on observed trajectories, interactions DDPM NGSIM 2023 ITSC ——
Yao et al. [94] vehicle trajectory prediction conditioned on observed trajectories, map DDPM Argoverse2 2023 CSIS-IAC ——
Diffuser [35] behavior planning unconditional;
@@
conditioned on rewards, constraints, skills;
@@
ADM D4RL 2023 ICLR ——
MPD [96] motion planning unconditional;
@@
Diffusion-ES [98] motion planning unconditional truncated DDPM nuPlan 2024 CVPR diffusion-es
@@
conditioned on adjacent views VLDM nuScenes 2024 CVPR Drive-WM
@@
conditioned on past frame, text VLDM WOMD et al. 2024 CVPR DriveAGI
CTG [36] vehicle trajectory generation conditioned on observed scene;
@@
ADM nuScenes 2023 ICRA CTG
CTG++ [111] multi-agent trajectory generation conditioned on observed scene;
@@
ADM nuScenes 2023 CoRL CTG++
Dragtraffic [51] multi-agent trajectory generation conditioned on initial scene, text LED WOMD 2024 IROS Dragtraffic
DJINN [47] multi-agent trajectory generation conditioned on arbitrary state;
classifier-free guidance;
@@
Argoverse2 2023 NIPS ——
Rempe et al. [114] human trajectory generation conditioned on observed scene;
@@
FDM [121] image-based driving scenario generation conditioned on previously sampled frames FDM Carla 2022 NIPS ——
GAIA-1 [38] image-based driving scenario generation conditioned on past image, text, action tokens;
@@
DriveDreamer-2 [122] image-based driving scenario generation conditioned on structured info by LLMs, text EDM nuScenes 2024 arxiv DriveDreamer2
@@
nuScenes 2024 CVPR panacea
@@
nuScenes 2023 arxiv DrivingDiffusion
WoVoGen [55] image-based driving scenario generation conditioned on past world volumes, actions, text, 2D image feature LDM nuScenes 2023 arxiv WoVoGen
LiDMs [56] point cloud-based driving scenario generation unconditional;
@@
Copilot4D [125] point cloud-based driving scenario generation conditioned on past observations, actions;
@@
DiffTraj [43] GPS trajectory generation conditioned on trip region, departure time;
@@
DiffSTG [23] traffic flow forecasting conditioned on past graph signals, graph structure NCSN PEMS et al. 2023 GIS DiffSTG
@@
conditioned on past graph signals feature, adjacency matrix DDPM PEMS et al. 2024 arxiv SpecSTG
@@
ST-SSPD [140] traffic flow forecasting conditioned on past data points, temporal encoding, node identifier DDPM METR-LA et al. 2023 MobiArch ——
@@
DiffTAD [147] trajectory anomaly detection unconditional DDIM NGSIM 2024 KBS ——
VAD [148] video anomaly detection unconditional;
@@
LDM, DDIM CUHK Avenue et al. 2023 ICCV ——